### v2

# Machine learning for hydrodynamic stability

Abstract.
A machine-learning strategy for investigating the stability of fluid flow problems is proposed herein. The objective is to exploit shallow neural networks for a binary classification task to distinguish between bifurcated and non-bifurcated states. The goal is to provide a simple yet robust methodology to find a nonlinear mapping from the parametric space to a bifurcating indicator identifying the region of interest. The computational procedure is demonstrably robust and does not require parameter tuning. The essential feature of the strategy is that the computational solution of the Navier–Stokes equations is a reliable proxy for laboratory experiments investigating sensitivity to flow parameters. The applicability of our bifurcation detection strategy is demonstrated by an investigation of two classical examples of flow instability associated with thermal convection. The codes used to generate and process the labelled data are available on GitHub.
Acknowledgements. This work was supported by EPSRC grants EP/W033801/1 and EP/Y028783/1.
1. Introduction
The issue of hydrodynamic stability has been extensively studied over the last century. Excellent books on the topic include Joseph [14], Landau & Lifshitz [15], Drazin & Reid [6] and Schmid & Henningson [18]. The graduate textbook by Drazin [5] provides a perfect introduction to the subject. Linearizations for small perturbations of a base (steady) flow are the starting point for developing mathematical theory and often provide results that match with laboratory experiments. Classical linear stability analysis assumes exponential behaviour of infinitesimal perturbations and leads to characteristic eigenvalue problems that can be analysed theoretically (following in the footsteps of Lord Rayleigh). They can also be explored computationally by discretising the governing partial differential equations (PDEs) linearised about the base flow and solving a high-dimensional eigenvalue problem to track the critical eigenvalues. See Cliffe et al. [4] for a comprehensive review and Wubs & Dijkstra [23, Chap. 6] for a contemporary treatment.
The concept of “learning from data” has become mainstream over the past decade with activity in all areas of science. As noted by Strang [21] the central goal of machine learning is to construct a function that classifies the training data correctly, so it can generalise to unseen test data. Assessing the effectiveness of a rudimentary (shallow) neural network in correctly classifying hydrodynamic stability is our primary focus here.
Starting from the pioneering paper of Raissi et al. [17], problems from fluid dynamics have featured prominently in papers exploring machine learning of solutions of partial differential equations. Efforts to train physics-informed neural networks (PINNs) in the context of fluid flow modelling have met with mixed success to date however. Wang et al. [22] identify the Stokes and Navier–Stokes equations as challenging problems that highlight some of the most prominent difficulties in training PINNs. Recent work of Pichi et al. [16] builds on earlier work by Chen et al. [3] in constructing a machine-learning strategy for learning the steady-state bifurcation structure of flow problems. The developed strategy is limited by a computationally expensive offline stage whereby a reduced basis set is computed from a singular-value decomposition of a (high-dimensional) snapshot matrix. The associated error in this dimensional reduction is difficult to quantify and limits the achievable accuracy when modelling fluid flow (because the decay of the singular values is so slow). In contrast, the methodology developed herein is based on direct simulation of the flow problem, where the effect of discretisation error can be monitored to ensure that it does not influence the resulting flow classification. Neural network training is compared with Gaussian process regression in a recent paper by Sousedik et al. [20] assessing the linear stability of two benchmark flow problems. The issues associated with the computation and careful tracking of eigenvalues are eschewed in our bifurcation detection strategy.
Our stability classification strategy involves two components. First and foremost, we need an accurate discretisation strategy for the governing PDEs modelling the time-dependent evolution of the flow. The second component is to specify a bifurcation detection strategy and assign a tolerance for labelling the training data. Precision is not so critical in this however—our experience is that the resulting neural network classification results are relatively insensitive to the choice of tolerance. A description of the particular machine-learning strategy that we employ can be found in section 2.
Regarding the discretisation strategy, a simple method-of-lines approach that combines a stable finite element spatial approximation together with a similarly accurate implicit time discretisation is adopted herein. One feature distinguishing our strategy from that of others is the use of self-adaptive time stepping that is designed to “follow the physics”. The benefit of this is that instability of the computed flow does not need to be explicitly excited. In singular cases (multiple steady-state flows) the onset of instability is triggered by computer round-off much as it would be triggered in physical experiments by geometrical imperfections in the laboratory apparatus. The upshot however, is that stable linear solver strategies are essential if we are to respect the singularity or ill-conditioning of the linearised algebraic equation systems. These issues are explored in the context of flow problem with a symmetry-breaking pitchfork flow bifurcation in section 3. The applicability of our bifurcation detection strategy is demonstrated in section 4, where we present results for two examples of flow instability associated with thermal convection that have been partially studied previously.
2. Bifurcation classification using a neural network
The two distinct flow problems considered in this work are defined on a bounded spatial domain $\Omega\in\mathbb{R}^{2}$ and a temporal domain $\tau=(0,\infty)$. Given a parabolic PDE operator $L$ (typically representing the Navier–Stokes equations) and a vector $\lambda\in\mathbb{R}^{n}$ of parameters, the goal is to explore the bifurcation structure of flow solutions $u({\vec{x}},t;\lambda)$ satisfying
$L(u({\vec{x}},t;\lambda))&=0&&\text{in $\Omega\times\tau$},\\ u(\cdot,t;\lambda)&=g(t;\lambda)&&\text{on $\partial\Omega\times\tau$},\end{aligned}$ (1)
starting from a quiescent flow profile at time $t=0$.
The characterisation of the bifurcation structure of a given flow problem is a two-stage procedure. In the first stage, approximations $u_{h}$ of solutions $u$ will be computed by discretising (1) in space (using mixed finite element approximation) and in time (using adaptive time stepping driven by local error control), by sampling parameter vectors $\lambda$ in the neighbourhood of the bifurcation boundary. [^1] This step requires a priori knowledge of the bifurcation phenomenon. While this can often be guessed by physical considerations or from laboratory experiments, determining suitable sample points would pose a challenge in cases where the system behaviour is totally unknown. A classification test is then applied to the sampled data to determine whether the computed flow is representative of the base flow or else has undergone a bifurcation. The result of this is a set of labelled data
[…]
where the label $\ell^{j}$ is either 0 or 1. A distinguishing feature of our time stepping strategy is that a bifurcated solution $(u_{h}^{j},1)$ typically evolves from the associated base state (this depends on the underlying bifurcation structure). Small perturbations of the flow solution are not explicitly introduced in the time stepping process. This supports our assertion that the computer can be viewed as a proxy for a corresponding laboratory experiment investigating sensitivity to flow parameters.
The second stage of our procedure is the generation of a surrogate model of the bifurcation structure. We will show that this can be done effectively by training a shallow neural network to classify the data. The machine-learning ingredients will be elementary: a shallow network will prove to be perfectly adequate and we do not need to carefully tune any hyperparameters. A random division of the data in $SS\SS$ into a training dataset and a test dataset will demonstrate that the machine-learned models of bifurcation structure are statistically valid. A representative surrogate generated with fixed PDE discretisation parameters can also be validated a posteriori, by sampling parameters close to the bifurcation boundary and comparing the model predictions with the labelled data that is generated for the same parameter values using a refined finite element mesh or a smaller time accuracy tolerance.
The neural network architecture was fixed throughout this study. The input to the network is the parameter vector $\lambda$ (two-dimensional in all cases discussed later) and the final output is a prediction of the corresponding label $\ell$. The network has a single hidden layer consisting of 32 neurons. Changing the width to 16 or to 64 neurons is not going to significantly affect the results. The activation function is the standard sigmoid. The output from the hidden layer is a two-dimensional vector $[z_{1},z_{2}]$ that is mapped onto a vector of probabilities $p=[p_{1},p_{2}]$ using a numerically stable softmax function
$z_{*}=\max\{z_{1},z_{2}\},\qquad p_{j}={e^{z_{j}-z_{*}}\over e^{z_{1}-z_{*}}+e^{z_{2}-z_{*}}},\quad j=1,2.$ (3)
The resulting output vector will be interpreted as the probability of computing a bifurcated solution ($p_{1}$) or the base solution ($p_{2}$). The training of the neural network is accomplished by comparing the output vector $p$ with the labelled output vector
$y=\begin{cases}[0,1]&\hbox{ if }\ell^{j}=0,\\ [1,0]&\hbox{ if }\ell^{j}=1.\end{cases}$ (4)
using a cross-entropy loss function (logistic regression) with the loss functional
$\mathcal{L}(y,p)=-y_{1}\ln(p_{1})-y_{2}\ln(p_{2}).$ (5)
In all cases discussed in later sections, the training of the network proved to be robust with respect to the initialisation of the weights and biases and did not require bespoke parameter tuning. [^2] The raw parameter data does need to be normalised however. This was done by subtracting the mean and dividing by the standard deviation of the individual components. A typical training run with ${m}=500$ data points with a fixed learning rate $\alpha=0.15$ is shown in Figure 1. The optimisation in this instance was 5000 passes through the dataset, taking a gradient descent step after every data point with gradients computed analytically by back-propagation. Looking more closely at the convergence history, we see that the output of the loss functional decreases non-monotonically initially with a reduction of an order of magnitude in the first 12 steps followed by a monotonic reduction by an additional factor of 2 (50%) after 36 steps. The rate of decrease of the loss functional is relatively slow thereafter: the relative reduction achieved in the final 4000 steps is close to 25%. The $\ell_{2}$ norm of the gradient vector when the training run ended was $1.49\times 10^{-7}$. The time taken for the complete training run was less than 4 minutes when run using the open-source Python code on a 2018 MacBook Pro.
Figure 1. Network training history for 500 data points (symmetric channel flow bifurcation).
The use of a neural network surrogate to model the stability of flow configurations with multiple parameters is completely novel as far as we are aware. The methodology will be shown to offer new insights into the assessment of the stability of the flow problems that feature in this study.
3. Validation of the classification strategy
The effectiveness of our classification strategy will be demonstrated in the first instance by studying a classical symmetry-breaking channel flow test problem and comparing results with those in the literature. Specifically, we compute approximations to the velocity $\vec{u}=[u_{x},u_{y}]$ and pressure $p$ satisfying the Navier–Stokes equations in a rectangular duct with a symmetric expansion. Details of the domain and a typical steady-state solution (below the critical Reynolds number with an unperturbed inflow) are shown in Figure 2. The associated steady-state boundary conditions are
  • parabolic flow $\vec{u}(-1,y)=(1-4y^{2},0)$ at the inflow boundary $(-1,y),\,|y|\leq 0.5$
  • natural conditions $\nu\frac{\partial u_{x}}{\partial x}=p$, $\frac{\partial u_{y}}{\partial x}=0$ at the outflow boundary $(16,y),\,|y|\leq 1$
  • no-flow conditions $\vec{u}=0$ along the channel walls $(x,\pm 1),0\leq x\leq 16$;
    $(x,\pm 0.5),-1\leq x\leq 0$; $(0,y)$, $0.5\leq|y|\leq 1$.
The position of the outflow boundary is far enough downstream that the flow is fully developed. The key feature of the base solution shown in Figure 2 is that it is reflectionally symmetric with respect to the centerline $y=0$, i.e., the stream function $\psi$ satisfies $\psi(x,y)=-\psi(x,-y)$. It follows that for the velocity,
$u_{x}(x,y)=u_{x}(x,-y),\quad u_{y}(x,y)=-u_{y}(x,-y).$ (6)
To be consistent with previous work, we will characterise the flow Reynolds number $\mathcal{R}=LU/\nu$ in terms of the kinematic viscosity $\nu$, the maximum speed at the inflow $U$ and the height of the channel at the outflow $L$. The symmetry of the base solution can be confirmed by observing that the two eddies downstream of the step in Figure 2 have the same reattachment length.
Refer to caption11222216161616
Figure 2. Symmetric step domain and computed flow solution for $\mathcal{R}=212.8$.
[…]
The TR–AB2 integrator is run in fully nonlinear mode with a second-order linearisation step followed by two fixed-point iterations (a total of three ${n_{x}}\times{n_{x}}$ sparse matrix solves at each time step). The integrator is run for a fixed number ${n_{t}}$ of steps with the time accuracy tolerance set to 3e-6. When generating labelled data, ${n_{t}}$ is set to 1200.
Figure 3. Computed steady-state flow solution for $\mathcal{R}=222.2$.
The specific flow problem is referred to as a 1:2 expansion in the computational study reported by Drikakis [7] and it exhibits a pitchfork bifurcation if the Reynolds number is increased past a critical value $\mathcal{R}^{*}$ of approximately 220. At the critical value the rightmost eigenvalue of the linearised Jacobian matrix (see Elman & Silvester [10]) changes from a real negative number (indicating linear stability) to a real positive number (instability). Just above the critical Reynolds number there are three steady solutions: the symmetric flow solution is unstable and there are two stable asymmetric solutions with the two recirculating eddies having a different length. One such solution computed at a Reynolds number $\mathcal{R}>\mathcal{R}^{*}$ is shown in Figure 3. The second steady state is the mirror image: the top and bottom eddies in Figure 3 are inverted.
Figure 4. Time step history and mean vorticity evolution for $\mathcal{R}=222.2$. The vertical lines refer to specific time steps that are identified in the text.
The mean vorticity $\omega(t)$, or the average vertical velocity at the outflow ($\partial\Omega_{N}$)
$\omega(t)=\int_{\Omega}\nabla\times\vec{u}(\cdot,t)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rm{d}\Omega}=\int_{\partial\Omega_{N}}\!\!\!\!u_{y}(\cdot,t)\,\rm{d}s,$ (7)
provides a convenient way of assessing the degree of departure from the steady symmetric flow (for which $\omega$=0). The evolution of the time step when computing the flow solution in Figure 3 is shown in Figure 4. Following the physics and checking snapshots of the evolving flow solution one can identify four distinct phases in the time integration process:
  • The first phase is the propagation of the inlet pulse from the inflow to the channel expansion, that is, the first time unit of integration. This milestone is reached after 136 time steps, at which point the step size is 0.017 and the two downstream eddies are just starting to form.
  • The second stage is the propagation of the inlet pulse from the channel expansion to the outlet. This milestone is reached after a further 270 time steps. At this point in time a symmetric downstream eddy profile is fully established. Following this, the time step grows rapidly (by two orders of magnitude) as the flow solution relaxes—an (unstable) symmetric eddy profile (similar to that in Figure 2) can be observed at step number 600.
  • The third phase is the transition from the symmetric eddy profile to an asymmetric profile. After 820 steps the mean vorticity rapidly increases by two orders of magnitude. An asymmetric eddy profile is clearly formed at step number 920.
  • The final phase is a relaxation into the final asymmetric eddy profile shown in Figure 3. The mean vorticity stops increasing at step 1020 before it settles to a steady value of $\omega=1.2\times 10^{-4}$. There is nothing to be gained by continuing the time integration beyond this point. Once the symmetry is broken there is no way back.
From a machine-learning perspective the mechanics of assigning a label to a computed flow solution needs to be definitive. The flow solution in Figure 2 is computationally symmetric because the $x$ coordinate where the upper eddy reattaches on the top boundary is close enough to the $x$ coordinate where the lower eddy reattaches on the bottom boundary. The asymmetry of the flow in Figure 3 is self-evident if we compare the computed vorticity (wall shear stress) on the top and bottom boundaries, as shown in Figure 5. The location $x>0$ where the wall shear stress is zero is the point where the eddy reattaches to the wall. These locations are obviously different: the vorticity on the lower boundary changes sign between boundary grid nodes 65 and 66, whereas the vorticity on the upper boundary changes sign between grid nodes 88 and 89. This simple test classifies the flow in Figure 3 as “bifurcated” ($\ell^{j}=1$). Applying the same test to the flow in Figure 2 the indices where the sign changes are within one grid point of each other—this classifies the flow as being symmetric ($\ell^{j}=0$). This is a very strict test of the flow symmetry—it is appropriate in cases where the loss of stability is a smooth function of the bifurcation parameter(s) and gives reliable results whenever the spatial resolution is sufficiently fine. [^4] If the sign change on the top was between nodes 74 and 75, and on the bottom was between nodes 73 and 74 or between 75 and 76 we would still classify the flow as being symmetric ($\ell^{j}=0$). Further justification will be provided later.
Figure 5. Computed wall shear stress for $\mathcal{R}=222.2$.
An issue that it is important to appreciate is that classical eigenvalue analysis provides no information about the size of asymmetric perturbation needed to the initiate a transition to an asymmetric flow profile from an unstable symmetric flow pattern. Herein, to replicate a practical laboratory experiment, the influence of perturbation magnitude will be studied. Specifically, we aim to assess the symmetry or otherwise of an evolving flow generated with a distorted steady-state parabolic inlet profile,
${u_{x}(-1,y)=\delta\cdot\sin(2\pi y)+(1-4y^{2})},$ (8)
where the parameter $\delta$ measures the size of the perturbation. The skewed inlet profile effectively disconnects the pitchfork bifurcation structure—the only asymmetric solution that is computed when $\delta$ is greater than unit roundoff ($\varepsilon$) is a flow profile with a longer eddy at the bottom compared to the top.
We generate data for two different sets of parameters $\lambda=(\mathcal{R},\delta)$. The first range of results is generated with 35 uniformly spaced Reynolds numbers between 209.6 and 225.7 and 10 logarithmically spaced perturbations between $10^{-16}$ and $10^{-3}$. The second range is generated with 30 uniformly spaced Reynolds numbers between 168.1 and 222.2 and 5 logarithmically spaced perturbations between $10^{-5}$ and $5\times 10^{-2}$. Combining these two data sets into a single set $SS\SS$, 281 data points were labelled as symmetric flows and 219 were classed as asymmetric.
Figure 6. Symmetric step channel flow: typical test data classification (left) and prediction accuracy (right). Red dots on the right are correct predictions, blue dots are incorrect.
Figure 7. Symmetric step channel flow: data classification (left); surrogate predictions from trained neural network (right). Red dots correspond to $\ell^{j}$=0 and blue dots correspond to $\ell^{j}$=1.
To test the predictive capability of the surrogate model, the data set $SS\SS$ is randomly divided into a training set with 375 data points and a test set with 125 data points. The classification of the test data in one such division is shown in Figure 6. In this representative data split the surrogate model generated by the training data correctly classifies 123 of the test data points with 2 asymmetric flows incorrectly classified as being symmetric. The test was repeated multiple times. The average precision after 5 independent tests was 0.964, and the average F1 score was 0.969. The visualisation of the bifurcation boundary in Figure 7 is generated by training the neural network on 375 data points, sampling the parameter set at 50,000 points normally distributed around $\mathcal{R}=218$ and plotting the predicted classification at these points.
Figure 8. Symmetric step channel flow: refined grid classification (left) and prediction accuracy (right). Red dots on the right are correct predictions, blue dots are incorrect.
The classification boundary in Figure 7 is consistent with estimates of the critical Reynolds number $\mathcal{R}_{*}$ for unperturbed flow ($\delta=0$) that have been published in the literature. Results for three studies reported in Table II in [7] suggest that $\mathcal{R}_{*}$ is between 216 and 219, but the authors also remark that “the exact value of the critical Reynolds number is difficult to fix because it depends on grid resolution, but it is close to 216.” This remark is supported by the eigenvalue results in [10, Fig. 4] showing that computed estimates of the critical eigenvalue governing linear stability move to the left as the spatial resolution is increased. To assess the sensitivity to discretisation parameters and validate our surrogate model, a second set of 50 classification data points were randomly generated close to the classification boundary with a refined spatial grid (level 6 in ifiss). In these calculations, the dimension ${n_{x}}$ of the algebraic system is increased from 23,842 to 94,146 and ${n_{t}}$, the number of time steps, is increased from 1200 to 1600. The results are shown in Figure 8. The only two parameter combinations that are misclassified by the coarse grid surrogate model have a perturbation of magnitude close to $10^{-4}$. Looking carefully at the results in the left plot in Figure 7 one can see a lot of uncertainty in the classification of flows with perturbation of this magnitude. We did not investigate this further. We note however that if the magnitude of the perturbation is increased to $5\times 10^{-3}$ then the classification results in Figure 7 suggest that the flow is not symmetric for all Reynolds numbers in the range considered.
4. Classification of instabilities in thermal convection
The broad applicability of our classification strategy will be demonstrated by exploring the bifurcation structure of two representative problems that arise when modelling buoyancy driven flow. The first case we consider is a symmetry-breaking pitchfork bifurcation and the second is a Hopf bifurcation from a steady flow solution to a periodic flow. We note that both problems are members of the set of canonical bifurcation test problems discussed by Wubs & Dijkstra in [23, Chap. 10].
4.1. A Rayleigh–Bénard convection problem
Following their discovery by Bénard in the laboratory in 1900, the development of recirculating convection cells has been extensively studied: both theoretically, see [5, Chap. 6], and computationally, see Gelfgat [12]. A state-of the-art algorithm that combines Newton’s method with a simple deflation strategy can be found in Boulle et al. [2].
The focus here will be on a geometric configuration that is outside the realm of Rayleigh’s shallow layer linearised stability analysis. Specifically we consider a (tall) 1:2 aspect ratio cavity with insulated vertical boundaries $x=0$ and $x=1$. The lower boundary $y=0$ is hot and the top boundary $y=2$ is cold. If the imposed temperature difference is sufficiently small then there is no motion and the buoyancy effect of rising fluid is balanced by the effect of gravity acting downward. If the temperature difference is increased beyond a critical value then the symmetry of the solution is broken and a steady recirculation is established that moves in a clockwise (upper branch) or an anticlockwise (lower branch) direction.
Figure 9. Computed isotherms for unperturbed Rayleigh–Bénard convection: steady conducting state (left) and bifurcated solution (right).
The fluid modelled is water so the Prandtl number Pr = $\nu/\epsilon$, where $\nu$ is the kinematic viscosity and $\epsilon$ is the thermal diffusivity, is set to 7.1. As in the previous section, we set up the bifurcation problem $\lambda$ having two parameters, Ra and $\delta$. The first of the parameters is the Rayleigh number that is associated with the standard nondimensionalization of the governing (Boussinesq) system, defined here to be
${\hbox{{Ra}}={g\beta(T_{\hbox{hot}}-T_{\hbox{cold}})d^{3}\over\nu\epsilon},}$
where $g$ is the acceleration due to gravity, $\beta$ is the thermal expansion coefficient of the fluid and $d$ is the distance between the hot and cold boundaries. The Rayleigh number conveniently characterises the degree of instability of the system. Computed steady-state temperature profiles for two values are illustrated in Figure 9. We observe that when Ra is 1510 the no-flow conducting state solution is stable, whereas when Ra is increased to 1530 it is unstable. The asymmetric solution shown in the right plot corresponds to an anticlockwise circulation. The other stable steady state when $\hbox{{Ra}}=1530$ features an clockwise circulation of the same magnitude with the isotherms reflected in the line of symmetry $x=0.5$. Note that, in this case, the kinetic energy
$k(t)={1\over 2}\int_{\Omega}\vec{u}(\cdot,t)\cdot\vec{u}(\cdot,t)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rm{d}\Omega},$ (9)
provides a convenient way of assessing the degree of departure from the equilibrium flow.
The second parameter within the vector $\lambda$ represents the magnitude of an asymmetric perturbation of the constant temperature at the base of the cavity. In addition, to model the action of “heating the pan” in a laboratory setting, the temperature variation on the horizontal walls is asymmetrically ramped up in time as was done in section 3. The resulting boundary conditions are
[…]
Figure 10. Step size history (left) and kinetic energy (right) for ${Ra}=1530$, $\delta=0$. The vertical lines refer to specific time steps that are highlighted in the text.
The time step evolution when computing the bifurcated solution in Figure 9 is shown in Figure 10. Following the physics and checking snapshots of the evolving temperature solution one can again identify four distinct phases in the time integration process:
  • The first phase is the development of the equilibrium solution as the horizontal walls are heated/cooled from zero to their final values of ±0.5plus-or-minus0.5$\pm$ 0.5± 0.5.555The visible jumps in the time step after 30 and after 270 steps correspond to an TR–AB2 averaging step. If averaging is not done then the time step stagnates before the equilibrium solution is established. Averaging was done every 30 steps until step 300, but was not invoked for the final 500 steps. A tiny jump in the kinetic energy between time steps 33 and 66 (too small to feature on the plot) is associated with the motion of the fluid needed to initiate the transition to the equilibrium state. A linear temperature profile is clearly established after 20 time units and 210 time steps.
  • After this milestone, the step size grows rapidly (by twelve orders of magnitude) as the temperature solution relaxes—an (unstable) symmetric temperature profile (as in the left plot in Figure 9) can be observed at time step 320.
  • The third phase is the transition from the symmetric profile to an asymmetric profile. After 330 steps the adaptive integrator starts rejecting steps and the step size is reduced by eleven orders of magnitude over the next 40 time steps. This decrease in step size is associated with a rapidly increase in the kinetic energy. An asymmetric temperature profile corresponding to an anticlockwise recirculation is clearly formed at time step 400. In contrast, when generating the solution with Ra=1510 shown in Figure 9, there is no increase in the kinetic energy after 300 steps and no associated reduction in the step size.
  • The final phase is the relaxation to the temperature profile illustrated in the right plot in Figure 9. The kinetic energy “overshoots” before being quickly damped to a perfectly steady asymmetric profile. As noted in section 3, there is nothing to be gained by continuing the integration beyond this point.
Figure 11. Rayleigh–Bénard convection: data classification (left); surrogate predictions from trained neural network (right). Red dots correspond to $\ell^{j}$=0 and blue dots correspond to $\ell^{j}$=1.
The dataset $SS\SS$ was generated by taking 60 uniformly spaced Rayleigh numbers between 1300 and 1600 and 7 logarithmically spaced perturbations between $10^{-16}$ and $10^{-2}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if the final kinetic energy was greater than ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\times 10^{-7}}$ (this is clearly the case for the parameter combination in Figure 10). With this test in place, 254 of the data points $SS\SS$ were labelled as a system in the conducting state and 166 were classed as asymmetric. The precise classification is shown in Figure 11. Sampling the trained neural network at 200,000 points leads to the prediction of the bifurcation boundary shown in the right plot. From this picture one would predict that the critical Rayleigh number for perfect uniform heating of the cavity bottom is in the interval $(1510,1520)$. The results also suggest that the critical Rayleigh number will be less than 1500 if the applied heating is skewed by a measurable amount $\geq 10^{-3}$, one way or the other. The results also indicate that if the perturbation is larger than $\geq 10^{-2}$, then the critical Rayleigh number will be less than 1300.
4.2. A differentially heated cavity problem
The second problem is motivated by laboratory experiments originally performed by Elder [8]. The experiments were performed over 50 years ago by observing the motion of aluminium powder suspended in viscous oil with a Prandtl number 1000. The aim here is to extend our understanding with a view to assessing the stability of the laboratory configuration for a variety of fluids with Pr ranging from 7.1 (water) to 1000 (glycerol). We note that the bifurcation structure for fluids with low Prandtl numbers has also been extensively studied over the past 25 years, see Gelfgat et al. [13]. Elder’s original experimental set up can be modelled as a (tall) 0.051:1 aspect ratio cavity with insulated horizontal boundaries $y=0$ and $y=1$. The left boundary $x=0$ is hot and the right boundary $x=0.051$ is cold. If the imposed temperature difference is sufficiently small then a steady flow solution will be established with a clockwise primary flow recirculation.
[…]
If the Rayleigh number is above a critical value then the steady solution loses stability via a Hopf bifurcation. The associated Rayleigh number that was identified from the laboratory experiments is $2.26\times 10^{9}\pm 30\%$ [^7] The definition of the Rayleigh number in [8] is scaled by a factor of ${d}^{3}$., the large uncertainty is attributed to the difficulty of detecting the onset of a very weak secondary flow. A precise critical value of $2.88\times 10^{9}$ is identified in [23, Sect. 10.4] by carefully tracking eigenvalues. Computational results for a Rayleigh number just above this value typically evolve to a periodic flow solution with small “ripples” along the vertical centreline $y=0.5$. Such an instability is apparent in the volume averaged kinetic energy history for a Rayleigh number of $2.89\times 10^{9}$ that is presented in Figure 12, wherein a stable oscillation with a period of 683 time units can be clearly identified after the first 1000 steps of integration. In contrast, when the computed kinetic energy is plotted on the same scale for a slightly smaller Rayleigh number, say Ra = $2.87\times 10^{9}$, there are no oscillations visible after 1000 time steps, see Figure 13.
Figure 12. Evolution of the kinetic energy for a differentially heated cavity with Ra=$2.89\times 10^{9}$ and Pr=1000: time steps 120–2000 (left), time steps 1200–2000 (right).
Figure 13. Evolution of the kinetic energy for a differentially heated cavity with Ra=$2.87\times 10^{9}$ and Pr=1000: time steps 120–2000 (left), time steps 1200–2000 (right).
There are only two distinct phases in the evolution of the time step when using stabilised TR–AB2 to compute the bifurcated solution in Figure 12:
  • The first phase is the development of the primary recirculation as the vertical walls are heated/cooled from zero to their final values of ±$\pm 0.5$. There is no analytical solution in this case—even a small temperature difference will generate a nontrivial flow in the cavity. The primary recirculation is fully developed after 100 time steps, which translates to about 200 units of time.
  • The second phase is the development of the oscillating solution. After the first 350 steps the step size variation is what one would expect when following the physics of a periodic solution.
Figure 14. Unsteady convection: training data classification (left) and prediction accuracy (right). Blue dots on the right are correct predictions.
Figure 15. Unsteady convection: surrogate predictions from the trained neural network sampled in the neighbourhood of the classification boundary.
The parameters defining the bifurcation problem that is the focus here are the Rayleigh number (Ra) and the Prandtl number (Pr). The dataset $SS\SS$ was generated for 7 specific values of Pr; namely, 7.1, 71, 140, 195, 250, 500 and 1000. We computed results for different values of Ra  between $2.81\times 10^{9}$ and $3.09\times 10^{9}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if, when looking at the final 800 time steps of the computation, the ratio of the difference between the maximum and the minimum energy $k$ divided by the mean value of $k$ is greater than ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}6\times 10^{-4}}$ (as is the case for the parameter combination in Figure 12). With this test in place, 45 of the data points $SS\SS$ were labelled as a system that reached a steady state and 49 points were classed as being oscillatory. The precise classification is shown in Figure 14.
Sampling the trained neural network at 200,000 points leads to the prediction of the bifurcation boundary that is shown in Figure 15. The results illustrated have the following distinctive features:
  • Both the raw data for Pr=1000 and the bifurcation point predicted by the trained neural network are in close agreement with the results in [23].
  • When the Prandtl number is reduced from 1000 by an order of magnitude, the critical Rayleigh number increases, eventually reaching a maximum value of $3.05\times 10^{9}$ when Pr is close to 100.
  • A change in behaviour can be clearly observed in the refined prediction of the stability boundary that is visualised in the rightmost plot. The results suggest that if Prandtl number is reduced below 100 then the transition to unsteady flow as a function of the Rayleigh number occurs significantly earlier.
Confirmation of these observations clearly requires an adaptive strategy. One needs to refine the surrogate approximation by generating additional data points in regions of parameter space to reduce the interpolation error where the bifurcation boundary is most uncertain. The design of a semi-rigorous adaptive strategy—where the interpolation error is balanced by the PDE approximation error—is left as future work.
5. Concluding remarks
While reliably trained neural networks offer the possibility of transformative technology in particular areas of science, a convincing case for replacing physics-based modelling by contemporary machine-learning strategies is still to be established at this time. The strategy presented herein has a very different flavour: the interpolating potential of a trained neural network is utilised in a role where it is known to be very effective.
Our use of a robust adaptive time-stepping strategy to follow the evolution of a flow from a quiescent state to a final steady or periodic state is a key feature of our strategy. Our impression is that integration of the Navier–Stokes equations in time has invariably been done by researchers in the past using fixed time steps. While fixed time-stepping methods can be effective in modelling the growth or otherwise of perturbations of the equilibrium flow, it cannot be used to model a laboratory experiment from start to finish.
While we have concentrated on tracking the first bifurcation point in this work, there is no reason why the methodology cannot be to adapted to locate a succession of branch points. All the flow problems considered in this work are representative of problems where a classical linearised stability analysis is predictive of reality in the sense of being able to reproduce results observed in laboratory experiments. One interesting future direction would be to consider applying the methodology to flow problems where transient growth of instabilities limits the applicability of classical eigenvalue analysis.
Reproducibility statement
The numerical results in this paper are all reproducible. The flow stability data was generated by running IFISS version 3.7 (https://www.manchester.ac.uk/ifiss) in batch mode. The IFISS software and user guide can be downloaded from the GitHub repository (https://github.com/mcbssds/IFISS_download). IFISS can be run using MATLAB (licensed by The MathWorks) or using the latest release of GNU Octave (available from https://octave.org). The data was processed using open-source software written in Python. The functions used to generate the figures in the paper and the MATLAB codes used to generate the batch files can all be downloaded from the GitHub paper repository: https://github.com/mcbssds/hydrostabilityNN.